In the Deacon process for the manufacture of chlorine, HCl and O2 react to form Cl2 and H2O. Sufficient

air (21 mole% O2, 79 mole% N2) is fed to provide 35% excess oxygen and the fractional conversion of

HCl is 85%. Calculate the composition (% mole) of the product stream components. (Basis: 100 moles

HCl)


1
Expert's answer
2021-04-16T05:10:26-0400

Basis : 100 moles of HCl

Fractional conversion of HCl = 85%

HCl reacted = 85 moles

HCl unreacted = 15 moles

Oxygen is supplied in 35% excess


"2HCl+\dfrac{1}{2}O_2\longrightarrow Cl_2+H_2O"


From Balanced chemical reaction,

2 moles of HCl reacts to form 1 mole of Cl2 and 1 mole of H2O

From 85 moles of HCl,

Cl2 formed = 42.5 moles

H2O formed = 42.5 moles

2 moles of HCl reacts with "\dfrac{1}{2}" moles of O2

85 moles of HCl reacts with "=85\times\dfrac{1}{4}=21.25" moles of O2

Moles of O2 supplied = "(\dfrac{35}{100}\times21.25)+21.25=28.68" moles

Moles of O2 unreacted = moles of O2 supplied "-" moles of O2 reacted = 7.43 moles

100 moles of air contains 21 moles of O2

28.68 moles of O2 is contained in 136.57 moles of air

100 moles of air contains 79 moles of N2

136.57 moles of air contains = 107.89 moles of N2


Product stream contains,

N2 = 107.89 moles = "\dfrac{107.89}{215.32}\times100=50.10\%"


O2 = 7.43 moles = "\dfrac{7.43}{215.32}\times100=3.45\%"


HCl = 15 moles = "\dfrac{15}{215.32}\times100=6.96\%"


Cl2 = 42.5 moles = "\dfrac{42.5}{215.32}\times100=19.73\%"


H2O = 42.5 moles = "\dfrac{42.5}{215.32}\times100=19.73\%"